Please Tutor help me solve this question. Find n if p(n,4) = 42 p(n,1)
Answer by ikleyn(52786)

.
p(n,4) = n*(n-1)*(n-2*(n-3).

p(n,1) = n.


Therefore, your equation is 

n*(n-1)*(n-2*(n-3) = 42*n.


Or, equivalently, 

(n-1)*(n-2)*(n-3) = 42.


It HAS NO solution in integer numbers.